INAF queue observing, 2017 Nov 25-26

Observer: F. Cusano, A. Gargiulo
Telescope Operator: S. Allanson
Support Astronomer: B. Rothberg
Instrument: LUCI, MODS, LBC

Summary

This night is shared with TSIP and we end observations at 7:06 UT.
At the sunset the sky is clear.
LUCI1 has still problems so we can not use it, the ad-sec dx is still under checking and can not be used for science so we can not use MODS2 and LUCI2. We start with MODS1.
Last minute communication (00:40) we can use the ad-sec dx for seeing limited observations but not for AO. We keep observing with MODS, we are already configured for it.
At 05:40 the TO has to close the dome for smoke and dust outside. From 9:30 the telescope is available again for INAF because TSIP night was a transit and they lost the right moment.

Night duration (12\xB0 el twilight) 1:40am- 7:06am UT = 8.5 h

Observing time: * *7h 20(%)

Weather time loss: * *h 30min (%)

technical time loss * *15min (%)

Data Summary

Project_RUN Objects Instrument FilterSorted ascending ExpTime Notes
LUCI Calibration Dark LUCI2      
MOS Calibration Flat and arc MODS      
Standard Feige 110 MODS dual grating    
Fabrizio Tri II MODS dual grating 2.5h x2 (bino) completed
Grazian I COSMOS308 MODS dual grating 1.75h x2 (bino) 4.5 on 7 obs
Caccianiga GB6J040159 +384927 MODS red grating 0.5h x2 (bino) completed
Caccianiga GB6J061110 +721814 MODS red grating 0.22h x2 (bino) 1 of 3

Details

00:57 slewing to Feige 110
  • acqBinoMODS mods1r.20171126.0002-3 ; mods2r.20171126.0002-3
  • execBinoMODS, dual grating mods1r.20171126.0004-6, mods1b.20171126.0002-4 ; mods2r.20171126.0004-6, mods2b.20171126.0002-4

01:18 slewing to Fabrizio, Tri II
  • acqBinoMODS Shallw1_592936.acq Shallw2_551353.obs, mods1r.20171126.0007-9 ; mods2r.20171126.0007-9
  • seeing 0.8" on guider at 1:25

01:35 execBinoMODS Shallw1_592936.obs Shallw2_551353.obs
  • mods1r.20171126.0010, mods1b.20171126.0005 ; mods2r.20171126.0010, mods2b.20171126.0005
  • seeing 0.75" on guider at 1:38

02:13 acqBinoMODS Deep1_551409.acq Deep2_551423.acq
  • aborted at 2:14, the TO checks the pointing

02:17 sending again acqBinoMODS Deep1_551409.acq Deep2_551423.acq
  • mods1r.20171126.0011-13 ; mods2r.20171126.0011-13
  • seeing 0.6" on guider at 2:19

02:29 execBinoMODS Deep1_551409.obs Deep2_551423.obs
  • mods1r.20171126.0014-17, mods1b.20171126.0006-9 ; mods2r.20171126.0014-17, mods2b.20171126.0006-9
  • seeing 0.7" on guider at 2:33
  • seeing 0.7" on guider at 3:08
  • seeing 0.6" on guider at 4:37

04:40 slewing to Caccianiga GB6J040159 +384927
  • problem in configuring MODS2
  • sending again acquisition at 4:47, it is working, acqBinoMODS GB6J040159 +384927.acq, mods1r.20171126.0019-21 ; mods2r.20171126.0018-20

04:57 execBinoMODS GB6J040159 +384927.obs
  • mods1r.20171126.0022-24 ; mods2r.20171126.0021-23
  • seeing 0.7" on guider at 4:58

05:36 switch to LUCI2

05:40 the TO has to close the dome for smoke and dust from outside

06:00 LUCI2 dark acquisition
  • Vanzella - DIT = 30; NDIT = 2 luci2.20171126.0001-0005
  • Vanzella SL - DIT = 2.51; NDIT = 20 luci2.20171126.0006-0010
  • Antoniucci - DIT = 2.51; NDIT = 24 luci2.20171126.00011-0015
  • Antoniucci - DIT 2.51; NDIT = 2 luci2.20171126.00016-0020

06:30 MODS mos calibration
  • MODS1 mask 592936 - MODS2 mask 551353
  • science flat mods1r.20171126.0025-26, mods1b.20171126.00010-11 ; mods2r.20171126.0024-25, mods2b.20171126.0010-11
  • lamp mods1r.20171126.0027-29, mods1b.20171126.00012-14 ; mods2r.20171126.0026-28, mods2b.20171126.0012-14

  • MODS1 mask 551409 - MODS2 mask 551423
  • science flat mods1r.20171126.0030-31, mods1b.20171126.00015-16 ; mods2r.20171126.0029-30, mods2b.20171126.0015-16
  • lamp mods1r.20171126.0032-24, mods1b.20171126.00017-19 ; mods2r.20171126.0031-33, mods2b.20171126.0017-19

09:24 we have the telescope again because the TSIP night was a transit time constrained and they lost the moment, we start with MODS

09:35 slewing to Caccianiga GB6J061110 +721814
  • acqBinoMODS mods1r.20171126.0034-36 ; mods2r.20171126.0035-37

09:49 execBinoMODS mods1r.20171126.0037 ; mods2r.20171126.0038
  • seeing 1" on guider at 9:50

10:07 slewing to Grazian I
  • acqBinoMODS pointI_mos.acq
  • preset cancelled,
  • 10:12 sending again pointI_mos.acq, mods1r.20171126.0038-40 ; mods2r.20171126.0039-41

10:24 execBinoMODS pointI_mos_1.obs
  • mods1r.20171126.0042-43, mods1b.20171126.0020-21 ; mods2r.20171126.0041-42, mods2b.20171126.0020-21
  • seeing 0.9" on guider at 10:27

11:00 execBinoMODS pointI_mos_2.obs
  • seeing 0.8" on guider at 11:01
  • mods1r.20171126.0044-45, mods1b.20171126.0022-23 ; mods2r.20171126.0043-44, mods2b.20171126.0022-23
  • we notice that the updatepointing in the script is causing the targets to drift away from the slit, we let the script go, but note that sequence for the first 4 spectra will be A,B,C,D; we corrected the other scripts to have ABAB again

11:42 execBinoMODS pointI_mos_3.obs
  • seeing 0.6" on guider at 11:51
  • mods1r.20171126.0046-47, mods1b.20171126.0024-25 ; mods2r.20171126.0045-46, mods2b.20171126.0024-25

12:19 execBinoMODS pointI_mos_4.obs
  • mods1r.20171126.0048, mods1b.20171126.0026 ; mods2r.20171126.0047, mods2b.20171126.0026
  • seeing 0.6" on guider at 12:21
  • we stop the ob at 12:39 because the last spectra will be in twilight

12:40 slew to Feige 34
  • preset cancelled, the TO check the pointing
  • 12:46 acqBinoMODS mods1r.20171126.0049-50 ; mods2r.20171126.0048-49

12:50 execBinoMODS Feige34
  • dual grating mods1r.20171126.0051-53, mods1b.20171126.0027-29 ; mods2r.20171126.0050-52, mods2b.20171126.0027-29

13:10 starting MODS calibrations for Grazian
  • flat mods1r.20171126.0054-55, mods1b.20171126.0030-31 ; mods2r.20171126.0053-54, mods2b.20171126.0030-31
  • arc mods1r.20171126.0056-58, mods1b.20171126.0032-34 ; mods2r.20171126.0055-57, mods2b.20171126.0032-34
Topic revision: r1 - 31 Jul 2024, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback